# Through the Magnifying Glass: Adaptive Perception Magnification for Hallucination-Free VLM Decoding

### v2

@@ 3.3 Attention-Based Magnification
4 Experiments and Results
Object-level Attribute-level MME Score POPE Accuracy
Existence Count Position Color Total* AOKVQA GQA COCO AVG*
Greedy 195.00 143.33 128.33 163.33 630.00 84.10 84.39 85.29 84.59
OPERA [24] 195.00 143.33 128.33 163.33 630.00 84.09 84.43 85.29 84.60
VCD [34] 183.33 143.33 110.00 150.00 586.67 84.18 82.32 85.49 83.99
ICD [57] 190.00 148.33 123.33 153.33 615.00 84.23 84.89 84.92 84.68
SID [25] 195.00 155.00 126.67 160.00 636.67 84.39 84.17 85.81 84.79
VDD [79] 190.00 140.00 133.33 165.00 628.33 86.73 85.52 86.71 86.32
PAI [43] 190.00 155.00 133.33 170.00 648.33 85.32 83.43 86.66 85.14
IBD [84] 190.00 160.00 133.33 170.00 653.33 86.37 84.80 86.65 85.94
API-C [73] 190.00 150.00 128.33 165.00 633.33 86.87 85.04 87.31 86.41
API-L [73] 190.00 150.00 133.33 168.33 641.67 86.52 84.30 86.60 85.81
ViCrop-R [76] 190.00 163.33 105.00 175.00 633.33 85.53 85.19 87.43 86.05
ViCrop-G [76] 195.00 153.33 96.67 175.00 620.00 86.09 85.18 87.31 86.19
PM (Ours) 195.00 175.00 138.33 175.00 683.33 86.94 85.46 87.68 86.70
Table 1: Evaluation results for MME Perception Scores (left) and POPE Accuracy (right). Columns marked with * indicate the overall scores for the benchmark. The result for BEAM is omitted for its equivalence with OPERA on single-token QA. API and ViCrop each propose multiple variants, denoted by postfixes (e.g., -C or -L), with details provided in the Supplementary Material. The proposed PM method consistently demonstrates superior performance on both benchmarks.
4.1 Experimental Settings
@@
We evaluate our method on three benchmarks: MME [20], POPE [37], and LLaVA-Bench [40], all commonly used for assessing multimodal hallucinations. For comparison, we include various decoding strategies, including greedy decoding, contrastive decoding-based methods (VCD [34], VDD [79], IBD [84], PAI [43], ICD [62], SID [25]), beam search [57], OPERA [24], API [73], and ViCrop [76]. Further details are provided in the Supplementary Material.
@@
We conduct our experiments using LLaVA-1.5 7B [41], a widely adopted VLM known for its strong visual interpretation and language understanding capabilities, as well as its computational efficiency. For hyperparameters, we start aggregating attentions from layer $\mathcal{L}=12$, set the scaling coefficient and kernel size to $\alpha=10$, $k=3$, and set the attention threshold for iterative refinement to $\beta=0.3$.
4.2 Quantitative Analysis
Tab. 1 presents the evaluation results of various methods on both the MME and POPE benchmarks. Our proposed method consistently outperforms all existing approaches by a substantial margin. Compared to MME, POPE primarily focuses on existence-based questions, allowing most methods to improve performance by reinforcing the contribution of visual tokens in model inference, thereby enhancing perception ability. In contrast, even within the perception subset, MME poses greater challenges by incorporating more complex questions that require both fine-grained visual recognition and reasoning.
Empirically, CD-based approaches operate at the logit level by contrasting visually intact or compromised outputs. While this strategy can enhance perception by suppressing hallucinated content, it fails to resolve cases where fine-grained details remain unrecognized across both logits. Methods that enhance visual representation at the embedding level, such as PAI and IBD, are less affected by this limitation, as they reinforce visual contributions at multiple layers. This enhancement may implicitly emphasize fine-grained details during intermediate processing, but it can also disrupt the balance between visual and linguistic reasoning, leading to potential trade-offs. By modifying input images, API strengthens visual contributions without altering the embedding distribution of VLM, but it primarily mitigates the impact of less relevant regions by applying a grayscale mask. While this reduces interference from irrelevant areas, it does not explicitly refine fine-grained details, limiting its enhancement on the perception of VLMs on subtle visual cues. ViCrop also exhibits limitations on MME, as isolating a region through cropping removes its surrounding visual context, making position-related questions more challenging. Additionally, providing the cropped image as an additional input can introduce ambiguity, hindering the VLM’s ability to interpret spatial relationships and maintain a coherent understanding of the scene.
In contrast, our method enhances critical visual information at the input level while preserving the VLM’s internal state, maintaining a stable embedding distribution across layers. It further improves fine-grained recognition by adaptively magnifying relevant regions while preserving contextual information. As a result, our approach effectively reduces hallucinations and consistently outperforms existing methods on benchmarks.
4.3 GPT-4o Assisted Evaluation
LLaVA-Bench COCO In-the-Wild
Greedy 81.15 92.69 95.76 56.94 71.83 79.95
BEAM 83.17 94.75 91.60 58.94 74.71 78.25
OPERA [24] 82.99 96.12 93.61 58.44 71.81 77.92
VCD [34] 82.00 92.21 94.39 57.17 69.4 81.22
ICD [57] 83.29 93.91 95.46 54.43 66.89 74.81
SID [25] 81.10 97.35 95.80 59.49 70.93 81.63
VDD [79] 81.38 95.62 102.75 61.60 78.22 83.01
PAI [43] 79.04 92.06 90.59 52.29 69.74 72.49
IBD [84] 85.29 95.73 95.45 61.59 77.91 82.86
PM (Ours) 85.76 97.74 95.80 62.46 77.21 84.50
Table 2: Results on LLaVA-Bench (%). Ref denotes reference-based evaluation, while Acc and Det represent accuracy and detailedness in vision-based evaluation, respectively.
Previous research [24, 84] has shown that GPT-4-assisted evaluation is a reliable metric for assessing hallucination. Accordingly, we adopt both reference-based and vision-based evaluations on LLaVA-Bench, using GPT-4o as the evaluator. In reference-based evaluation, GPT-4o receives the question with a detailed textual description of the image, while in vision-based evaluation, it is provided with both the question and the image.
Under both settings, GPT-4o scores the answer generated by the tested method and a reference answer from GPT-4 on a scale of 0 to 10. The final score is computed as the ratio of the tested method’s score to GPT-4’s score. For reference-based evaluation, GPT-4o focuses on overall content quality, whereas for vision-based evaluation, it examines the image and assesses accuracy and detailedness. We exclude API [73] and ViCrop [76] from these benchmarks, as they are not designed for open-end generation. Complete evaluation prompts are provided in the Supplementary.
4.4 Maintaining Cognition Ability
MME-Cognition CS NR Trans CR Total*
Greedy 112.86 70.00 107.50 67.50 357.86
BEAM 112.86 47.50 90.00 60.00 310.36
OPERA [24] 112.86 47.50 90.00 60.00 310.36
VCD [34] 120.00 37.50 97.50 72.50 327.50
ICD [62] 120.71 70.00 92.50 65.00 348.21
SID [25] 122.86 55.00 85.00 65.00 327.86
VDD [79] 110.14 58.00 68.00 86.00 322.64
PAI [43] 116.43 70.00 100.00 67.50 353.93
IBD [84] 117.14 70.00 107.50 57.50 352.14
API-C [73] 115.00 47.50 107.50 47.50 317.50
API-L [73] 120.00 45.00 115.00 45.00 325.00
ViCrop-R [76] 115.00 50.00 105.00 65.00 335.00
ViCrop-G [76] 109.29 50.00 105.00 86.00 350.29
PM (Ours) 117.86 42.50 107.50 87.50 355.36
To investigate this, we evaluate our method and baselines on the Cognition split of MME, where answering requires reasoning beyond direct visual cues. As shown in Tab. 3, most CD-based methods that explicitly modify output logits exhibit a notable decline in cognition ability, aligning with expectations. Surprisingly, PAI and IBD, which enhance the attention weight of visual tokens through intermediate embedding modifications, do not incur a substantial loss in cognition performance. This suggests that the bias introduced by their layer-wise manipulation is relatively mild and can be compensated for by subsequent model layers. Furthermore, despite API-based methods modifies the image without altering the VLM’s internal reasoning, they still lead to degraded performance. We hypothesize that the gray masking applied to images in the cognition set where backgrounds are typically uniform may introduce additional perceptual burden, ultimately impacting overall performance.
4.5 Qualitative Analysis
@@
Figure 3: Qualitative examples on the POPE benchmark showing the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
Figure 3 presents qualitative examples illustrating the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In the first example, the VLM initially attends to the top-left region, recognizing the potential presence of a chair but failing to identify it. With the magnified input, the chair is presented at a higher spatial resolution, enabling the VLM to correctly recognize it. Similarly, in another case, the VLM misidentifies an object behind a door window as a tree. However, the perception-magnified input allows the model to re-examine the object with greater detail, preventing hallucinated predictions.
@@
4.6 Ablations
MME-Perception Object-level Attribute-level
Existence count position color Total
@@ 4.6 Ablations
PM 195.00 175.00 138.33 175.00 683.33
Table 4: Ablations on iterative refinement (IR) and multi-layer aggregation (MLA). Both components contribute to performance improvements compared to their respective baselines.
@@
We ablate the impact of iterative refinement in constructing the perception map for the magnifier. As shown in Tab. 4, the refined perception maps consistently outperforms their counterpart without iterative refinement, demonstrating the effectiveness of our approach. These results indicate that iterative refinement alleviates the effects of information aggregation during perception map construction, leading to more precise coverage of visually relevant regions for next-token generation. While iterative refinement substantially enhances the accuracy of the perception map, it does not entirely prevent information leakage from masked tokens. This limitation arises from the inherent feature mixing among visual tokens during early-stage processing, as most VLMs employ a Vision Transformer [16] for feature extraction. Despite this, our method achieves comprehensive coverage of key visual regions while maintaining a strong balance between attention precision and computational efficiency, making it a robust and effective solution for guiding the magnifier.
MME-Perception Object-level Attribute-level
Existence Count Position Color Total
Greedy 195.00 143.33 128.33 163.33 630.00
API CLIP [73] 190.00 155.00 131.67 170.00 646.67
API LLaVA [73] 190.00 165.00 126.67 170.00 651.67
ViCrop-R [76] 195.00 163.33 116.67 170.00 645.00
ViCrop-G [76] 195.00 158.33 116.67 170.00 640.00
Rel Map [4] 195.00 175.00 130.00 155.00 655.00
PM w/o IR 190.00 175.00 135.00 165.00 665.00
Table 5: Comparison of different perception maps for guiding the magnifier. IR denotes iterative refinement. Our perception map construction method achieves superior results even without IR.
MME-Perception Object-level Attribute-level
Existence count position color Total
Greedy 195.00 143.33 128.33 163.33 630.00
@@
Crop 195.00 155.00 128.33 170.00 648.33
Crop & Resize 195.00 153.33 123.33 175.00 646.67
Magnification 195.00 175.00 138.33 175.00 683.33
Table 6: Comparison of methods for highlighting key visual regions based on the perception map.
@@
We also ablate the impact of multi-layer aggregation in constructing the perception map for the magnifier. As shown in Tab. 4, perception maps generated by aggregating attention patterns across multiple layers exhibit substantial differences from those constructed using attention from a single layer. Empirically, while VLMs tend to focus more on visual features in middle-to-deep layers, the specific layer receiving the strongest visual attention varies unpredictably. This observation suggests that relying on a single layer may lead to inconsistencies in capturing visual information. Our analysis shows that in most cases, a random subset of intermediate layers exhibits a clear and structured visual attention distribution, whereas in other layers, the model allocates little to no attention to visual tokens, apart from a few sink tokens [66]. This variability highlights the necessity of aggregating attention patterns across multiple layers to ensure that the most attended visual features across layers are effectively incorporated into the perception map.
4.7 Quality of Perception Map
Existing studies [4, 55] use attention-based relevancy maps to improve VLM interpretability by aggregating attention weights with backpropagated gradients. API [73] constructs attribution maps using either CLIP’s CLS token decomposition or averaged attention weights in LLaVA, while ViCrop [76] introduces attention aggregation techniques that vary in their reliance on gradient-based information. Although developed for different purposes, these methods can also be adapted for perception map construction.
Building on these methods, we compare their heatmap quality against our perception map, using the base version from Sec. 3.1 without iterative refinement and applying identical post-processing. As shown in Tab. 5, our approach achieves superior perception quality. CLIP-based maps prioritize global image-text similarity rather than token-level relevance, while API-LLaVA and ViCrop-R aggregate attention from a single intermediate layer and average across heads, leading to reduced stability and diluted key features. In contrast, our max-pooling approach better retains critical visual information. The Rel map method and ViCrop-G employ backpropagation-based fine-grained attributions but provide limited benefits for magnification, as they focus on token-level attribution with more precise attention values but without necessarily improving coverage. Furthermore, their reliance on computationally expensive gradient updates makes them less practical for efficient perception-guided magnification.
4.8 Image Prompting Method
@@
Various methods have been proposed to “prompt” images to emphasize specific regions [69, 76], including add a bounding box on the highlighted region, or apply blurring effect, crop the remaining regions, or crop and resize the key area. In this section, we benchmark these methods as alternatives to our magnifier, evaluating their effectiveness in highlighting key visual regions for focus-adaptive decoding when combined with the constructed perception map. As shown in Tab. 6, our magnification approach effectively leverages the perception map to enhance the perception ability of the VLM, whereas other prompting techniques provide only marginal improvements over greedy decoding.
@@
This difference arises because magnification not only highlights important regions but also reveals additional fine-grained details within limited resolution, allowing the model to capture subtle visual patterns. In contrast, many prompting methods merely emphasize region importance without increasing effective resolution, limiting their impact on fine-grained recognition. Meanwhile, when combined with ViCrop, the proposed method still improves performance, as ViCrop also enhances local details through cropping and resizing. However, ViCrop discards contextual information from the image, leading to reduced effectiveness in tasks that require spatial awareness, such as position-related questions.
5 Conclusion

### v3

@@ Methods
Object-level Attribute-level MME MME POPE Accuracy
Existence Count Position Color Percep* Cognition AOKVQA GQA COCO AVG*
Greedy 195.00 143.33 128.33 163.33 630.00 357.86 84.10 84.39 85.29 84.59
OPERA 195.00 143.33 128.33 163.33 630.00 310.36 84.09 84.43 85.29 84.60
VCD 190.00 143.33 120.00 155.00 608.33 327.50 84.47 83.13 85.61 84.40
ICD 190.00 148.33 123.33 153.33 615.00 348.21 84.23 84.89 84.92 84.68
SID 195.00 155.00 126.67 160.00 636.67 327.86 84.39 84.17 85.81 84.79
VDD 190.00 140.00 133.33 165.00 628.33 322.64 86.73 85.52 86.71 86.32
M3ID 190.00 150.00 133.33 166.67 640.00 339.36 86.52 85.46 86.73 86.24
PAI 190.00 155.00 133.33 170.00 648.33 353.93 85.32 83.43 86.66 85.14
IBD 190.00 160.00 133.33 170.00 653.33 352.14 86.37 84.80 86.65 85.94
API-C 190.00 150.00 128.33 165.00 633.33 317.50 86.87 85.04 87.31 86.41
API-L 190.00 150.00 133.33 168.33 641.67 325.00 86.52 84.30 86.60 85.81
ViCrop-R 190.00 163.33 105.00 175.00 633.33 335.00 85.53 85.19 87.43 86.05
ViCrop-G 195.00 153.33 96.67 175.00 620.00 350.29 86.09 85.18 87.31 86.19
[…]
Table 1: Evaluation results for MME Perception Scores (left), Cognition Scores (mid), and POPE Accuracy (right). Columns marked with * indicate the overall hallucination scores for the benchmark. The result for BEAM is omitted for its equivalence with OPERA on single-token QA. API and ViCrop each propose multiple variants, denoted by postfixes (e.g., -C or -L), with details provided in the Supplementary. The proposed PM method demonstrates superior performance on both benchmarks.
Experiments and Results
@@
We evaluate our method on three benchmarks: MME (Fu et al. 2023), POPE (Li et al. 2023c), and LLaVA-Bench (Liu et al. 2023), all commonly used for assessing multimodal hallucination. For comparison, we include various decoding or inference strategies, including greedy decoding, VCD (Leng et al. 2024), VDD (Zhang et al. 2024), M3ID (Favero et al. 2024), IBD (Zhu et al. 2024), PAI (Liu, Zheng, and Chen 2025), ICD (Wang et al. 2024a), SID (Huo et al. 2024), beam search (Sutskever, Vinyals, and Le 2014), OPERA (Huang et al. 2024), API (Yu, Yu, and Wang 2024), and ViCrop (Zhang et al. 2025a). Further details are provided in the Supplementary.
@@
We conduct our experiments using LLaVA-1.5 7B (Liu et al. 2024b). We start aggregating attentions from layer $\mathcal{L}=12$, set the scaling coefficient and kernel size to $\alpha=10$, $k=3$, and set the attention threshold for iterative refinement to $\beta=0.3$.
Quantitative Analysis
Table 1 reports results on MME and POPE. POPE centers on existence-based queries, where most methods benefit from reinforcing visual token importance. MME is more challenging, with its perception subset targeting hallucination and its cognition subset requiring higher-level reasoning. While our method focuses on mitigating hallucination, we report cognition results for completeness, noting that most baselines underperform on this subset as modifications on the embedding and logit could harm the VLM reasoning.
Among baselines, CD-based methods suppress hallucinated logits but often fail when visual detail remains ambiguous across both logits. Embedding-level methods such as PAI and IBD improve perception by reinforcing visual features across multiple layers. However, they may still struggle when relevant regions are too small or diffuse to be effectively represented in the ViT backbone. API suppresses irrelevant regions via masking, but does not enhance salient visual content. ViCrop isolates key areas via cropping, which may hinder spatial understanding due to loss of context and potential ambiguity from dual-image inputs, resulting in limited MME gains. In contrast, our method magnifies the relevant regions while preserving surrounding context and internal states of the VLM. This preserves the embedding distribution and enhances fine-grained recognition without compromising spatial coherence or reasoning, leading to state-of-the-art results across both benchmarks.
In addition, all methods exhibit varying degrees of degradation on the MME cognition subset, whereas ours does not. A detailed analysis of how and why different methods affect the cognitive ability is provided in the Supplementary.
@@
Figure 3: Qualitative examples on the POPE benchmark showing the input question, the original image, the raw and processed perception map $\mathcal{H}^{*}$ and $\mathcal{P}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
GPT-4o Assisted Evaluation
LLaVA-Bench COCO In-the-Wild
Greedy 81.15 92.69 95.76 56.94 71.83 79.95
BEAM 83.17 94.75 91.60 58.94 74.71 78.25
OPERA 82.99 96.12 93.61 58.44 71.81 77.92
VCD 82.00 92.21 94.39 57.17 69.4 81.22
ICD 83.29 93.91 95.46 54.43 66.89 74.81
SID 81.10 97.35 95.80 59.49 70.93 81.63
VDD 81.38 95.62 102.75 61.60 78.22 83.01
M3ID 83.02 96.41 95.78 61.88 77.49 82.83
PAI 79.04 92.06 90.59 52.29 69.74 72.49
IBD 85.29 95.73 95.45 61.59 77.91 82.86
[…]
Table 2: Results on LLaVA-Bench (%). Ref denotes reference aggrement, while Acc and Det represent accuracy and detailedness in vision-based evaluation.
Previous studies (Huang et al. 2024; Zhu et al. 2024) have demonstrated the reliability of GPT-4-assisted evaluation for hallucination assessment. We extend their evaluation protocol and experiment on LLaVA-Bench to benchmark the open-end generation quality. Overall, GPT-4o scores responses based on reference agreement (Ref), visual accuracy (Acc), and descriptive detail (Det), depending on the evaluation setting. We omit API and ViCrop as they are not applicable to open-ended generation. Details on the GPT-4o metrics and prompts can be found in the Supplementary.
Qualitative Analysis
Figure 3 presents qualitative examples illustrating the input question, the original image, the raw and processed perception map $\mathcal{H}^{*}$ and $\mathcal{P}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In the first example, the VLM initially attends to the top-left region, recognizing the potential presence of a chair but failing to identify it. With the magnified input, the chair is presented at a higher spatial resolution, enabling the VLM to correctly recognize it. Similarly, in another case, the VLM misidentifies an object behind a door window as a tree. However, the perception-magnified input allows the model to re-examine the object with greater detail, preventing hallucinated predictions.
@@
Ablations
MME-Perception Object-level Attribute-level Percep
Existence count position color Total*
@@
[…]
Table 3: Ablations on iterative refinement (IR) and multi-layer aggregation (MLA). Both contribute to performance improvements compared to their respective baselines.
@@
We ablate the effect of iterative refinement in constructing the perception map. As shown in Tab. 3, the refined maps consistently outperform their non-iterative counterparts, confirming the benefit of this strategy in improving visual localization. Iterative refinement helps address tokens overlooked from the early-stage aggregation of visual features, resulting in more accurate coverage of relevant regions. Although some residual leakage from masked tokens remains due to information exchange in the ViT (Dosovitskiy et al. 2021) visual feature extractor, our approach still achieves decent grounding of key areas with minimal computational overhead, making it an effective and efficient solution for guiding the magnifier.
MME-Perception Object-level Attribute-level Percep
Existence Count Position Color Total*
(a) Perception Map Construction Methods
API CLIP 190.00 155.00 131.67 170.00 646.67
API LLaVA 190.00 165.00 126.67 170.00 651.67
ViCrop-R 195.00 163.33 116.67 170.00 645.00
ViCrop-G 195.00 158.33 116.67 170.00 640.00
Rel Map 195.00 175.00 130.00 155.00 655.00
[…]
(b) Perception-Guided Magnification Methods
@@
Masking 195.00 155.00 128.33 170.00 648.33
ViCrop 195.00 153.33 123.33 175.00 646.67
[…]
Table 4: Ablation results across two groups: (a) comparing different methods for constructing perception maps; (b) comparing prompting techniques that highlight key regions based on the map. IR denotes iterative refinement.
@@
We ablate the effect of aggregating attention across multiple layers when constructing the perception map. As shown in Tab. 3, multi-layer aggregation yields significantly better coverage of relevant visual regions compared to using attention from a single layer.
While middle-to-deep layers generally focus more on visual content, the most informative attention varies across layers and samples. In many cases, only a subset of intermediate layers exhibits structured visual attention, while others focus primarily on sink tokens (Xiao et al. 2024). Aggregating across layers reduces this variability, ensuring that key visual cues are reliably captured for guiding the magnifier.
Comparison with Different Maps.
Attention-based visualization methods (Chefer, Gur, and Wolf 2021; Stan et al. 2024) have been proposed to improve VLM interpretability by combining attention weights with gradients. We adapt several representative techniques—including CLIP-based Rel(evance) Maps, API (Yu, Yu, and Wang 2024), and ViCrop (Zhang et al. 2025a), to construct a map to guide the magnifier, and compare them to our approach w/o IR using identical post-processing. As shown in Tab. 4, our method achieves superior coverage of relevant visual regions. Attribution maps based on CLIP often reflect global similarity rather than token-level saliency, while API and ViCrop variants aggregate attention from a single layer, resulting in lower stability and diluted focus. Backpropagation-based methods (e.g., Rel Map, ViCrop-G) offer fine-grained attribution but contribute little to magnification quality and are computationally expensive. In contrast, the proposed strategy, even w/o the IR process, preserves critical visual cues with better spatial consistency and minimal overhead, making it well-suited for efficient perception-guided decoding.
Image Prompting Methods.
@@
Various methods have been proposed to “prompt” images to emphasize specific regions (Yang et al. 2024; Zhang et al. 2025a), including add a bounding box, apply blurring effect, masking out the remaining regions, or crop and resize the key area as additional input (ViCrop). We benchmark these methods as alternatives to our magnifier, evaluating their effectiveness in highlighting key visual regions when combined with the constructed perception map. As shown in Tab. 4, our magnification approach more effectively leverages the perception map to enhance the perception ability of the VLM.
@@
This difference arises because magnification not only highlights important regions but also reveals additional fine-grained details within limited resolution, allowing the model to capture subtle visual patterns. In contrast, many prompting methods merely emphasize region importance without increasing effective resolution, limiting their impact on fine-grained recognition. Meanwhile, ViCrop also yields certain improvement, as it also enhances local details with resized key regions. However, ViCrop discards contextual information and confuses the VLM by supplying the additional resized image, leading to reduced effectiveness in tasks such as position-related and counting questions.
[…]
Figure 4: Hyperparameter analysis of $\alpha$ and $\beta$ on MME. PM demonstrates robust performance across different settings.
Hyperparameter Analysis on $\alpha$ and $\beta$.
We analyze the effect of $\alpha$ and $\beta$, which control magnification strength and the stopping criterion for iterative refinement, respectively. Increasing $\alpha$ enhances heatmap contrast and strengthens magnification, while decreasing $\beta$ increases refinement steps, improving coverage but at higher computational cost (capped at 8 iterations). As shown in Fig. 4, smaller $\alpha$ pairs well with low $\beta$, which already ensures broad region coverage, reducing the need for strong magnification. Overall, the perception magnifier remains robust across a wide range of hyperparameters, consistently identifying key visual regions. Notably, using a larger $\beta$ approximates the performance of the basic PM w/o iterative refinement, yet substantially mitigates hallucinations with only $2\times$ the compute.
Conclusion